At 27°C the average total energy of O2 is approximately -
 
 
 

Grade:12th pass

1 Answers

Arun
4 years ago
Total kinetic energy = n (3/2 RT0
where n = Number of moles of the gas
R = Gas constant
T = Absolute temperature
Molecular weight of methane,
CH4 = 12 + 4 * 1 = 16
∴ Number of moles of methane in 8.0 gm of methane
= 8.0/16.0 = 0.5
R = 8.314 joules/K/mole
T = 27 + 273 = 300 K
∴ Total kinetic energy of the molecules in 8.0 gm of methane at 27oC = InI * 3/2 RT
= 0.5 * 3/2 * 8.314 * 300 = 1870.65 joules
∴ Average kinetic energy = 1870.65/6.023 *1023 *0.5
= 6.21 * 10-21 joules/molecule